Solve each of the following quadratic equations by factorising.
1/3x^2 - 1/2x = - 1/6

Answer by Alan3354(69443)

Solve each of the following quadratic equations by factorising.
1/3x^2 - 1/2x = - 1/6
---
Multiply by 6
2x^2 - 3x + 1 = 0
(2x - 1)*(x - 1) = 0